OpenStudy (anonymous):

simplify sin (360º-x)

Parth (parthkohli):

Sine is periodic every \(360^{\circ}\). This means that \(\sin(x) = \sin(x \pm 360^{\circ})\). Here, \(\sin(360^{\circ} - x ) = \sin(360^{\circ} - x - 360^{\circ}) = \sin(-x) = -\sin(x)\)